OpenStudy (anonymous):

To solve an equation in x, you simplify the equation so that it becomes easy to tell which value(s) of x make the equation true. A.simplify -6x + 13 + 7x = 6 + x + 7 until you have the equation in the form x = ?. Tell what value(s) of x make the equation true. B. What can you conclude about the original equation?

OpenStudy (anonymous):

\[-6x + 13 + 7x = 6 + x + 7\] combine like terms on the left and right which one do you want to do first?

OpenStudy (anonymous):

if it is not clear what i am asking, what is \(6+7\) ?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

ok so now we have \[-6x+13+7x=x+13\] now what is \(7x-6x\) ?

OpenStudy (anonymous):

13x ?

OpenStudy (anonymous):

oh no dear, you are subtracting not adding \[7x+6x=13x\] but you have \[7x-6x\]

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

Derp. 1x?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

btw if you don't want your math teacher to think you are not clever, don't write \(1x\) just write \(x\)

OpenStudy (anonymous):

that means we have \[x+13=x+13\] what can you say about the left and the right?

OpenStudy (anonymous):

they look the same to you?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

B. What can you conclude about the original equation? they are the same thing, so it is true for any value of \(x\)

OpenStudy (anonymous):

clear enough?

OpenStudy (anonymous):

Yea